# Sparse Voxels Rasterization: Real-time High-fidelity Radiance Field Rendering

### v1

@@ B.2 Dynamic Morton Order
where $f^{\mathrm{(k)}}:[0\cdots 7]\mapsto[0\cdots 7]$ is one of the eight permutation mappings. The pseudocode for computing the ray sign bits and the mapping function from Octree Morton code to dynamic Morton order is provided in Sec. B.1.
@@ B.3 Proof of Correct Ordering
“For all sparse voxel sets in $V^{\ell}$, their dynamic Morton order is always aligned with the near-to-far rendering order for all rays with (+,+,+)(+,+,+)( + , + , + ) direction signs.”
@@
When $\ell{=}1$, there is only one Octree level. The dynamic Morton order of the eight voxels for (+,+,+)(+,+,+)( + , + , + ) ray directions is illustrated in Fig. 10. The bit field from most to least significant bit is for $x$, $y$, and $z$ directions, respectively. As the ray is going toward $+x$ direction, we can always render the voxels in the $-x$ side ($000,001,010,011$) first before the voxels in the $+x$ side ($100,101,110,111$), which is aligned with the most significant bit of the Morton order. Similarly, for the voxels in the $-x$ side, we can render the voxels in the $-y$ side ($000,001$) before the $+y$ side ($010,011$) as the ray is going toward $+y$. Finally, we can see that the rendering order is correct if we iterate the voxels following the assigned Morton order for ray with (+,+,+)(+,+,+)( + , + , + ) directions.
@@
We want to show $S(\ell)\implies S(\ell+1)$ is true. For any sparse voxel set $w\in V^{\ell+1}$, there exists a sparse voxel set $v\in V^{\ell}$ that can evolve into $w$ by: i) selecting a subset of voxels in $v$ to subdivide with the source voxels removed and ii) removing some of the voxels. The $S(\ell)$ indicates that the dynamic Morton order of $v$ has the correct rendering order. To extend for a new Octree level, three zero bits are first append to the least significant bit of the Morton order of every voxel in $v$, which does not affect the ordering. When subdividing a voxel, the eight child voxels share the same most significant $3\ell$ bits as the source voxel, while the least significant $3$ bits follow the same dynamic Morton order as in the base case Fig. 10. This reflects the fact that the new child voxels should keep the same relative order to the other voxels as their source parent voxels as the child voxels are all in the 3D space of the source voxels. The rendering ordering of the eight child voxels can also follow the same Morton order as the base case. That is the Morton order is still rendering-order correct after subdividing some voxels in $v$. Finally, removing voxels does not affect the ordering of the remaining others. In sum, the Morton order of $w$ also has the correct rendering order so $S(\ell)$ implies $S(\ell+1)$. By induction, $S(\ell)$ is true for all positive integer $\ell$.
Figure 10: Base case. Dynamic Morton order for (+,+,+)(+,+,+)( + , + , + ) ray direction signs under the base case with 1 Octree level. The three bits from left to right is for the $x$, $y$, and $z$ directions respectively. The rendering order is correct for all rays going toward (+,+,+)(+,+,+)( + , + , + ) direction. See Sec. B.3 for more details.
@@ C Additional Implementation Details
We start the optimization from empty space with raw density set to $h_{\mathrm{geo}}{=}{-}10$. We use spherical harmonic (SH) with $N_{\mathrm{shd}}{=}3$ degrees. The learning rate is set to $0.05$ for the grid point densities, $0.01$ for zero-degree SH coefficients, and $0.00025$ for higher-degree SH coefficients. The momentum and the epsilon value of the Adam optimizer are set to $(0.1,0.99)$ and $1\text{e}{-}15$. The initial Octree level is $h_{\mathrm{lv}}{=}6$ (i.e., $64^{3}$ voxels) for the bounded scenes and the foreground main region of the unbounded scenes. To model unbounded scenes, we use $h_{\mathrm{out}}{=}5$ background shell levels with $h_{\mathrm{ratio}}{=}2$ times the number of foreground voxels. We use average frame color as the color coming from infinite far away for unbounded scenes. The early ray stopping threshold is set to $h_{\mathrm{T}}{=}1\mathrm{e}{-4}$ and the supersampling scale is set to $h_{\mathrm{ss}}{=}1.1$. Inside each voxel, we sample $K{=}1$ point for novel-view synthesis and $K{=}3$ points for the mesh reconstruction task.
We train our model for $20{,}000$ iterations. We find the benefit of training longer is marginal. The voxels are subdivided every $h_{\mathrm{every}}{=}1{,}000$ iterations until $15{,}000$ iterations, where the voxels with top $h_{\mathrm{percent}}{=}5$ percent priority are subdivided each time. We set $h_{\mathrm{rate}}{=}1$ and skip subdividing voxels with a maximum sampling rate below $2h_{\mathrm{rate}}$. The voxels are pruned every $h_{\mathrm{every}}{=}1{,}000$ iterations until $18{,}000$ iterations, where voxels with maximum blending weights less than a pruning threshold are removed. The pruning threshold is linearly increased from $0.0001$ at the first pruning to $h_{\mathrm{prune}}{=}0.05$ at the last pruning.
The loss weights are set to $\lambda_{\mathrm{ssim}}{=}0.02$, $\lambda_{\mathrm{T}}{=}0.01$, $\lambda_{\mathrm{dist}}{=}0.1$ after $10{,}000$ training iterations, $\lambda_{\mathrm{R}}{=}0.01$, $\lambda_{\mathrm{tv}}{=}1\text{e}{-}10$ until $10{,}000$ training iterations. For the mesh reconstruction task, the weights for normal-depth alignment self-consistency loss are set to $\lambda_{\mathrm{n\text{-}dmean}}{=}0.001$ and $\lambda_{\mathrm{n\text{-}dmed}}{=}0.001$ for mean and median depth respectively. The initial depths and normals are bad so the two normal-depth consistency loss is activated at the later training iterations. We find the median depth converges the fastest so we activate median depth-normal consistency loss at $3{,}000$ iterations, which also only regularizes the rendered depth as median depth is not differentiable. The mean depth-normal consistency loss is activated at $10{,}000$ iterations.
@@ E More Results
bonsai 215 147 18.3 12.5 0.204 0.178 31.89 31.47 0.942 0.937 1.2M 6.1M
counter 160 102 20.7 14.6 0.199 0.184 29.03 28.62 0.909 0.896 1.2M 7.3M
kitchen 128 97 25.0 13.1 0.126 0.119 31.47 31.42 0.927 0.927 1.8M 7.9M
room 153 146 21.1 13.6 0.218 0.190 31.44 30.89 0.919 0.917 1.5M 8.8M
@@
bicycle 72 164 31.9 10.8 0.211 0.206 25.18 24.93 0.765 0.755 6.1M 9.0M
garden 81 137 33.1 9.4 0.107 0.119 27.39 26.87 0.867 0.849 5.9M 8.7M
stump 110 142 25.5 11.2 0.216 0.219 26.61 26.11 0.772 0.756 4.9M 9.1M
treehill 123 170 22.4 11.8 0.327 0.284 22.47 22.24 0.632 0.609 3.7M 9.2M
flowers 137 125 22.0 12.1 0.335 0.277 21.57 21.40 0.606 0.617 3.6M 9.2M
@@
train 206 103 11.3 12.4 0.206 0.210 22.11 20.72 0.816 0.788 1.1M 9.3M
truck 154 117 16.3 10.3 0.147 0.119 25.40 24.74 0.882 0.873 2.6M 9.1M
@@
chair 418 225 5.4 3.3 0.012 0.014 35.89 35.84 0.987 0.986 0.3M 2.5M
drums 406 275 5.8 2.9 0.037 0.044 26.16 26.10 0.955 0.946 0.3M 2.0M
ficus 476 399 5.0 2.1 0.012 0.015 34.85 34.35 0.987 0.984 0.3M 1.2M
hotdog 596 233 5.2 3.6 0.020 0.022 37.67 37.44 0.985 0.983 0.1M 2.6M
lego 415 175 5.9 4.4 0.015 0.017 35.77 35.38 0.983 0.980 0.3M 3.8M
materials 575 225 5.3 3.4 0.034 0.040 30.01 29.92 0.960 0.953 0.3M 2.5M
mic 344 361 5.6 2.0 0.006 0.008 35.38 35.72 0.991 0.991 0.3M 1.0M
ship 254 107 7.9 7.8 0.107 0.110 30.92 30.34 0.907 0.883 0.3M 5.9M
Table 19: Per-scene side-by-side comparison of our base setup with 3DGS [19].
@@
Acknowledgements.
We thank Min-Hung Chen and Yueh-Hua Wu for helpful paper flow suggestions and proofreading.

### v3

@@ B.2 Direction-dependent Morton Order
where $f^{\mathrm{(k)}}:[0\cdots 7]\mapsto[0\cdots 7]$ is one of the eight permutation mappings. The pseudocode for computing the ray sign bits and the mapping function from Octree Morton code to direction-dependent Morton order is provided in Sec. B.1.
@@ B.3 Proof of Correct Ordering
“For all sparse voxel sets in $V^{\ell}$, their direction-dependent Morton order is always aligned with the near-to-far rendering order for all rays with (+,+,+)(+,+,+)( + , + , + ) direction signs.”
@@
When $\ell{=}1$, there is only one Octree level. The direction-dependent Morton order of the eight voxels for (+,+,+)(+,+,+)( + , + , + ) ray directions is illustrated in Fig. 10. The bit field from most to least significant bit is for $x$, $y$, and $z$ directions, respectively. As the ray is going toward $+x$ direction, we can always render the voxels in the $-x$ side ($000,001,010,011$) first before the voxels in the $+x$ side ($100,101,110,111$), which is aligned with the most significant bit of the Morton order. Similarly, for the voxels in the $-x$ side, we can render the voxels in the $-y$ side ($000,001$) before the $+y$ side ($010,011$) as the ray is going toward $+y$. Finally, we can see that the rendering order is correct if we iterate the voxels following the assigned Morton order for ray with (+,+,+)(+,+,+)( + , + , + ) directions.
@@
We want to show $S(\ell)\implies S(\ell+1)$ is true. For any sparse voxel set $w\in V^{\ell+1}$, there exists a sparse voxel set $v\in V^{\ell}$ that can evolve into $w$ by: i) selecting a subset of voxels in $v$ to subdivide with the source voxels removed and ii) removing some of the voxels. The $S(\ell)$ indicates that the direction-dependent Morton order of $v$ has the correct rendering order. To extend for a new Octree level, three zero bits are first append to the least significant bit of the Morton order of every voxel in $v$, which does not affect the ordering. When subdividing a voxel, the eight child voxels share the same most significant $3\ell$ bits as the source voxel, while the least significant $3$ bits follow the same direction-dependent Morton order as in the base case Fig. 10. This reflects the fact that the new child voxels should keep the same relative order to the other voxels as their source parent voxels as the child voxels are all in the 3D space of the source voxels. The rendering ordering of the eight child voxels can also follow the same Morton order as the base case. That is the Morton order is still rendering-order correct after subdividing some voxels in $v$. Finally, removing voxels does not affect the ordering of the remaining others. In sum, the Morton order of $w$ also has the correct rendering order so $S(\ell)$ implies $S(\ell+1)$. By induction, $S(\ell)$ is true for all positive integer $\ell$.
Figure 10: Base case. Direction-dependent Morton order for (+,+,+)(+,+,+)( + , + , + ) ray direction signs under the base case with 1 Octree level. The three bits from left to right is for the $x$, $y$, and $z$ directions respectively. The rendering order is correct for all rays going toward (+,+,+)(+,+,+)( + , + , + ) direction. See Sec. B.3 for more details.
@@ C Additional Implementation Details
We start the optimization from empty space with raw density set to $h_{\mathrm{geo}}{=}{-}10$. We use spherical harmonic (SH) with $N_{\mathrm{shd}}{=}3$ degrees. The learning rate is set to $0.025$ for the grid point densities, $0.01$ for zero-degree SH coefficients, and $0.00025$ for higher-degree SH coefficients. We decay all learning rates by $0.1$ at the 19K iteration. The momentum and the epsilon value of the Adam optimizer are set to $(0.1,0.99)$ and $1\text{e}{-}15$. The initial Octree level is $h_{\mathrm{lv}}{=}6$ (i.e., $64^{3}$ voxels) for the bounded scenes and the foreground main region of the unbounded scenes. To model unbounded scenes, we use $h_{\mathrm{out}}{=}5$ background shell levels with $h_{\mathrm{ratio}}{=}2$ times the number of foreground voxels. We use average frame color as the color coming from infinite far away for unbounded scenes. The early ray stopping threshold is set to $h_{\mathrm{T}}{=}1\mathrm{e}{-4}$ and the supersampling scale is set to $h_{\mathrm{ss}}{=}1.5$. Inside each voxel, we sample $K{=}1$ point for novel-view synthesis and $K{=}3$ points for the mesh reconstruction task.
We train our model for $20K$ iterations. The voxels are subdivided every $h_{\mathrm{every}}{=}1K$ iterations until $15K$ iterations, where the voxels with top $h_{\mathrm{percent}}{=}5$ percent priority are subdivided each time. We set $h_{\mathrm{rate}}{=}1$ and skip subdividing voxels with a maximum sampling rate below $2h_{\mathrm{rate}}$. The voxels are pruned every $h_{\mathrm{every}}{=}1K$ iterations until $18K$ iterations, where voxels with maximum blending weights less than a pruning threshold are removed. The pruning threshold is linearly increased from $0.0001$ at the first pruning to $h_{\mathrm{prune}}{=}0.05$ at the last pruning.
The loss weights are set to $\lambda_{\mathrm{ssim}}{=}0.02$, $\lambda_{\mathrm{T}}{=}0.01$, $\lambda_{\mathrm{dist}}{=}0.1$ after $10K$ iterations, $\lambda_{\mathrm{R}}{=}0.01$, $\lambda_{\mathrm{tv}}{=}1\text{e}{-}10$ until $10K$ iterations. For the mesh reconstruction task, the weights for normal-depth alignment self-consistency loss are set to $\lambda_{\mathrm{n\text{-}dmean}}{=}0.001$ and $\lambda_{\mathrm{n\text{-}dmed}}{=}0.001$ for mean and median depth respectively. The initial depths and normals are bad so the two normal-depth consistency loss is activated at the later training iterations. We find the median depth converges the fastest so we activate median depth-normal consistency loss at $3K$ iterations, which also only regularizes the rendered depth as median depth is not differentiable. The mean depth-normal consistency loss is activated at $10K$ iterations.
@@ E More Results
Comparison with popping-resistant 3DGS variants.
More comparisons with recent 3DGS variants are in Tab. 19. 3DGS [20] has popping artifacts due to the ordering and overlapping issues. StopThePop [39] uses running sort and 3DGRT [33] uses ray tracing for accurate ordering but drops FPS by 28% and 67% respectively. EVER [28] further handles the Gaussian overlapping cases but with even less FPS. Our method ensures correct ordering (Sec. 3.1.2) with FPS and quality comparable to the original 3DGS.
3DGS variants Ours
Method 3DGS [20] StopThePop [39] 3DGRT [33] EVER [28] fast-rend base
No popping $\triangle$ $\triangle$
[…]
† Re-evaluated on our machine using the public code.
Table 19: Comparison with 3DGS variants tackling popping artifact on Mip-NeRF360 dataset [2]. The LPIPS values here are evaluated with the correct intensity scale between $[-1,1]$ following EVER [28]. 3DGRT and EVER use ray tracing approach instead of rasterization. EVER solves the Gaussians ordering and overlapping issues but sacrificing more FPS.
Results on Scannet++.
Scannet++ [58] is a large-scale dataset covering various types of indoor scenes. To reconstruct the bounded indoor environments, we heuristically set the scene center as the camera centroid and the scene radius as twice the maximum camera distance from the centroid. The voxel grid starts at $64^{3}$ without a background region. Additionally, we implement ray density ascending regularization and a spherical harmonic reset trick, which we find improves results on the public validation set. The result on the held-out test-set is shown in Tab. 20. Our method achieves good results on all metrics. Some indoor fly-through videos are provided in the released code.
[…]
Small set (12 scenes)
Plenoxels [59] 0.399 22.177 0.841
TensoRF [5] 0.404 23.524 0.857
INGP [34] 0.363 23.695 0.871
Zip-NeRF [3] 0.320 24.630 0.887
Nerfacto [48] 0.340 23.498 0.868
3DGS [20] 0.312 23.389 0.876
FeatSplat [29] 0.303 24.177 0.880
RPBG [61] 0.271 24.005 0.882
Ours 0.300 24.365 0.886
Full set (50 scenes)
TensoRF [5] 0.406 24.022 0.850
Zip-NeRF [3] 0.325 25.041 0.880
3DGS [20] 0.319 23.893 0.871
Ours 0.313 24.709 0.874
Table 20: Scannet++ [58] indoor dataset. The test-set images are not released to prevent overfitting. We submit our rendering results to the scannet++ official website for a 3rd-party evaluation. The online benchmark is: https://kaldir.vc.in.tum.de/scannetpp/benchmark/nvs. In average, our training time is 12 minutes per scene; our rendering FPS is 197 at $1752\times 1168$ resolutions. Voxel size statistic is: 13.61% <3mm, 19.25% 3-5mm, 32.43% 5mm-1cm, 23.31% 1-2cm, 6.66% 2-3cm, 4.73% >3cm. We do not use the sparse points prior from COLMAP [43] in this submission.
@@
bonsai 215 128 18.3 15.3 0.204 0.171 31.89 31.51 0.942 0.944 1.2M 6.6M
counter 160 85 20.7 18.7 0.199 0.176 29.03 28.72 0.909 0.905 1.2M 8.4M
kitchen 128 78 25.0 17.8 0.126 0.112 31.47 31.29 0.927 0.934 1.8M 9.2M
room 153 131 21.1 17.1 0.218 0.185 31.44 31.10 0.919 0.924 1.5M 8.9M
@@
bicycle 72 147 31.9 13.5 0.211 0.190 25.18 25.29 0.765 0.773 6.1M 9.2M
garden 81 118 33.1 12.4 0.107 0.106 27.39 27.31 0.867 0.865 5.9M 9.6M
stump 110 129 25.5 13.0 0.216 0.206 26.61 26.38 0.772 0.769 4.9M 9.2M
treehill 123 157 22.4 13.7 0.327 0.262 22.47 22.74 0.632 0.646 3.7M 9.4M
flowers 137 120 22.0 14.4 0.335 0.268 21.57 21.72 0.606 0.637 3.6M 9.4M
@@
train 206 127 11.3 11.4 0.206 0.186 22.11 21.26 0.816 0.813 1.1M 8.3M
truck 154 129 16.3 11.0 0.147 0.100 25.40 25.00 0.882 0.888 2.6M 9.0M
Table 21: Real-world datasets for per-scene side-by-side comparison with 3DGS [20]. Our result here is the base setup. We show average results of our 2x faster rendering and 3x faster training variants in the main paper.
[…]
Scene 3DGS ours 3DGS ours 3DGS ours 3DGS ours 3DGS ours 3DGS ours
@@
chair 418 197 5.4 4.6 0.012 0.013 35.89 35.91 0.987 0.986 0.3M 3.0M
drums 406 241 5.8 4.1 0.037 0.043 26.16 26.09 0.955 0.947 0.3M 2.3M
ficus 476 360 5.0 3.1 0.012 0.014 34.85 34.37 0.987 0.984 0.3M 1.3M
hotdog 596 218 5.2 4.8 0.020 0.019 37.67 37.42 0.985 0.984 0.1M 2.8M
lego 415 156 5.9 5.8 0.015 0.016 35.77 35.54 0.983 0.981 0.3M 4.3M
materials 575 213 5.3 4.5 0.034 0.037 30.01 30.00 0.960 0.954 0.3M 2.7M
mic 344 328 5.6 3.0 0.006 0.007 35.38 36.00 0.991 0.992 0.3M 1.1M
ship 254 111 7.9 8.6 0.107 0.106 30.92 30.38 0.907 0.886 0.3M 5.7M
Table 22: Synthetic object-centric dataset for per-scene side-by-side comparison with 3DGS [20]. Our overall quality is slightly worse than 3DGS on this dataset while the synthetic object-centric scenario is off our main focus.